OpenStudy (anonymous):

how do you do y=1/8x-6 into standard form?

OpenStudy (anonymous):

standard form is AX+BY=C 1/8x-y=6

OpenStudy (timothywhitaker):

then you multipuly 1/8 and y so (8)(1/8x-y=(6)(8)

OpenStudy (anonymous):

Okay